### v1

# Higher order Jacobi method for solving a system of linear equations

Abstract
This work proposes a higher-order iterative framework for solving matrix equations, inspired by the structure and functionality of neural networks. A modification of the classical Jacobi method is introduced to compute higher-order coefficient matrices through matrix-matrix multiplications. The resulting method, termed the higher-order Jacobi method, structurally resembles a shallow linear network and allows direct computation of the inverse of the coefficient matrix. Building on this, an iterative scheme is developed that, once the network parameters are determined for a known system, enables efficient resolution of system variations without re-computing the coefficients. This iterative process naturally assumes the form of a deep recurrent neural network. The proposed approach moves beyond conventional Physics-Informed Neural Networks (PINNs) by providing an explicit, training-free definition of network parameters rooted in physical and mathematical formulations. Computational analysis demonstrates improved order of complexity, suggesting a promising direction for algorithmic efficiency in solving linear systems. This methodology opens avenues for interpretable and scalable solutions to physically motivated problems in computational science.
1 Introduction
Many real-world phenomena, ranging from fluid dynamics and heat transfer to financial modeling and electromagnetism, are governed by Partial Differential Equations (PDEs). These equations describe how physical quantities evolve over space and time. However, solving PDEs analytically is often impractical, necessitating numerical methods to calculate their solutions.
To enable numerical solutions, partial differential equations (PDEs) [1, 2, 3] are discretized, converting the continuous problem into a system of algebraic equations through finite-difference or finite-element approximations. This discretization process leads to a matrix equation of the form shown in Eq. (1).
$A\cdot x=b$ (1)
where $A$ is a system matrix derived from discretization, $x$ is the vector of unknowns, and $b$ represents known constraints or forcing terms. Efficiently solving this system is fundamental in computational science [4, 5], enabling simulations [6, 7, 8] and optimizations [9, 10] across various engineering and scientific domains.
There are several approaches to solving the matrix equation. Direct methods [11, 12, 13], such as Gaussian elimination [14] and LU decomposition [15], solve for $x$ exactly when feasible. Iterative methods [16, 17], such as the Jacobi [18, 19] and conjugate gradient methods [20, 21, 22], are effective for large, sparse systems. The AI driven approach, physics-informed neural networks (PINNs) [23, 24, 25], leverages deep learning to approximate solutions while respecting underlying physical laws.
This article presents a variation of the Jacobi method that can harness emerging hardware, such as GPUs being developed for AI to solve matrix equations. With this approach, higher-order system matrices are computed recursively, ultimately forming a deep linear network. The weights and biases are explicitly derived, offering deeper insight into PINNs, which are often treated as black boxes. It also eliminates the significant time and computational cost required for training. Notably, this higher-order method (HOM) can also be seamlessly extended to handle dense matrices, a scenario where traditional Jacobi methods are typically less effective.
Although the current state of physics-informed neural networks largely emphasizes physics-informed training [26, 27], our work shifts focus toward intuitive, physics-informed explicit generation of weights and biases, paving the way for a more interpretable framework where the physics itself defines the structure of the network.
This study demonstrates explicit weight and bias construction for linear systems. Future extensions to nonlinear systems, building upon the linear case as a foundation, can offer even deeper insights into PINNs, particularly where the governing physics defines the nonlinear transformation within the network.
2 The Higher Order Method
The Jacobi approach to solving a matrix equation follows a decomposition of the matrix $A$ into its diagonal $D$, lower $L$, and upper $U$ matrices. The solution is then iteratively computed until the solution converges within a desirable tolerance, as shown in Eq. (2).
$x^{(k+1)}=D^{-1}\cdot(b-(L+U)\cdot x^{(k)})$ (2)
where $k$ and $k+1$ represent the current and next iteration steps in the iterative process.
2.1 The higher order Jacobi method for solving matrix equation
To facilitate the application of the higher-order version of the Jacobi iteration method, the system matrix $A$ and the source vector $b$ are normalized with respect to the diagonal matrix $D$ as defined in Eq. (2). The resulting reformulated system is presented in Eq. (3).
$x^{(k+1)}=-(A_{n}-I)\cdot x^{(k)}+b_{n}$ (3)
where $I$ is the identity matrix. $A_{n}$ and $b_{n}$ are normalized matrices of $A$ and $b$, respectively.
Further simplification of Eq. (3) to express the next iteration as a linear combination of the current one leads to Eq. (4).
$x^{(k+1)}=\alpha\cdot x^{(k)}+\beta$ (4)
where
$\alpha=-(A_{n}-I)$ (5)
and
$\beta=b_{n}$ (6)
This transformation ensures a more structured and scalable computation.
Starting with an initial guess $x^{(0)}$, the solution after the first iteration $x^{(1)}$ is calculated as shown in Eq. (7)
$x^{(1)}=\alpha\cdot x^{(0)}+\beta$ (7)
Similarly, the second iteration is computed as shown in Eq. (8)
$x^{(2)}=\alpha\cdot x^{(1)}+\beta$ (8)
However, substituting Eq. (7) into Eq. (8), corresponding to the second-order approximation, enables direct computation of the second iterate from the initial guess, resulting in the following.
$x^{(2)}=\alpha\cdot(\alpha\cdot x^{(0)}+\beta)+\beta$ (9)
Expanding Eq. (9) results in Eq. (10),
$x^{(2)}=\alpha^{2}\cdot x^{(0)}+(\alpha+I)\cdot\beta$ (10)
The simplification of Eq. (10) yields Eq. (11), explicitly illustrating the linear dependency of the second iterate on the initial guess.
$x^{(2)}=\alpha^{(1)}\cdot x^{(0)}+\beta^{(1)}$ (11)
where
$\alpha^{(1)}=\alpha^{2}$ (12)
and
$\beta^{(1)}=(\alpha+I)\cdot\beta$ (13)
Further generalizing Eq. (11),
$x^{(2k)}=\alpha^{(1)}\cdot x^{(0)}+\beta^{(1)}$ (14)
which allows iterations to be performed using the second-order coefficients, $\alpha^{(1)}$ and $\beta^{(1)}$, respectively, instead of the first-order coefficients $\alpha$ and $\beta$. Consequently, a single iteration with the second-order coefficients is equivalent to two iterations with the first-order coefficients.
However, Eq. (11) takes the same form as Eq. (7), i.e., having the linear dependence on the initial guess. Following the same procedure of calculating second higher order coefficients, recursively substituting the higher order equation into itself, the general form of the higher order Jacobi method is represented in Eq. (15).
$x^{(2^{k})}=\alpha^{(k)}\cdot x^{(0)}+\beta^{(k)}$ (15)
where the recursive definitions for the higher-order coefficients are given by
$\alpha^{(k)}={(\alpha^{(k-1)})}^{2}$ (16)
and
$\beta^{(k)}=(\alpha^{(k-1)}+I)\cdot\beta^{(k-1)}$ (17)
Consequently, the solution after $2^{k}$ iterations with first-order coefficients is equivalent to 1 iteration with the $k^{th}$-order coefficients.
This exponential reduction in the number of iterations comes at the cost of increased computational complexity. Each higher-order step requires multiplying two square matrices, whereas traditional Jacobi iterations only involve multiplying a square matrix with a column vector.
For large matrices, block matrix multiplication and GPU acceleration with parallel processing significantly reduce real-world execution time of computing the higher-order coefficients.
2.2 Finding inverse of coefficient matrix from the higher order Jacobi method
The expanded form of Eq. (15) exhibits a clear structural similarity to a deep linear network comprising $k$ layers, with analytically defined weights and biases at each layer. It is well-established that any linear network, irrespective of its depth, can be equivalently represented as a single-layer network. In the context of the proposed higher-order Jacobi method, this single-layer equivalent corresponds to the inverse of the normalized coefficient matrix $A_{n}$, as elaborated in the following discussion.
In Eq. (15), the solution uniquely depends on the bias term $\beta^{(k-1)}$, while the other term is arbitrary, even though $\alpha^{(k-1)}$ is uniquely defined, as the initial guess can be arbitrary. The term $\alpha$ times $x$ can become insignificant with null vector as an initial guess. Thus, the multiplication term can be omitted, resulting in Eq. (18).
$x^{(2^{k})}=\beta^{(k)}$ (18)
Expanding Eq. (18) recursively results in Eq. (19).
$x^{(2^{k})}=(\alpha^{(k-1)}+I)\cdot(\alpha^{(k-2)}+I)\quad\ldots\quad(\alpha^{(1)}+I)\cdot(\alpha+I)\cdot\beta^{(0)}$ (19)
where
$\beta^{(0)}=\beta$ (20)
Expressing Eq. (19) using compact product notation leads to the form shown in Eq. (21).
$x^{(2^{k})}=\left(\prod_{i=k-1}^{0}(\alpha^{(i)}+I)\right)\cdot\beta$ (21)
where
$\alpha^{(0)}=\alpha$ (22)
Abstracting Eq. (21), results in Eq. (23).
$x^{(2^{k})}=A_{n}^{-1}\cdot\beta$ (23)
where
$A_{n}^{-1}=\prod_{i=k-1}^{0}(\alpha^{(i)}+I)$ (24)
The explicit closed-form representation of the inverse of matrix $A_{n}$, as shown in Eq. (24), is derived from an iterative Jacobi method, which proves the equivalence of direct and iterative methods of solving the matrix equation, whenever necessary conditions for convergence are met in the regular Jacobi iteration method [28, 29, 30]. In addition, the order of accuracy can be selected; the higher the order of coefficient matrices, the higher the accuracy. Eq. (23) can be renormalized to obtain the actual inverse of the original system matrix $A$.
2.3 The higher order Jacobi method as a deep network
Although Eq. (15) takes the same form as a deep linear network with weights and biases, Eq. (21) represents its transformation into a single-layer linear network, especially one without biases.
For an arbitrary forcing matrix $b$, the higher-order method produces a solution similar to the regular Jacobi iteration method. However, in real-world applications, the system matrix $A$ can be altered due to the material properties embedded in different PDEs [31] or due to mesh refinement [32], which requires the recalculation of higher-order matrices, which is computationally expensive.
This recalculation, even if there is a single change in one value of the original coefficient matrix, has the same computational complexity as matrix-matrix multiplication, making it inefficient for dynamically changing systems, especially for non-linear systems, where coefficients depend on the solution. This is where an iterative approach finds its way in and gives insight into importance/need of layers in neural networks. One such process is discussed in the following section.
2.4 Induced source iterative scheme as an explicit recurrent neural network
While the coefficient matrix may change dramatically in some of the situations such as mesh refinement, the following scheme is proposed for slowly varying coefficient matrix such as in nonlinear systems and even eigenvalue problems. Instead of recalculating the coefficients, the goal is to modify the source matrix such that solving the modified source matrix [33, 34] with a known inverse produces the same result as solving the desired system without recalculating its inverse. This approach eliminates the need for recomputing higher-order coefficients while maintaining solution consistency.
Let the modified source be $b_{m}$ as shown in Eq. (25).
$b_{m}=b+b_{i}$ (25)
where $b$ is the original source matrix for which the new system will be solved and $b_{i}$ is the induced source that will be computed iteratively.
This iterative approach introduces the following series of intermediate steps in which the solution is first obtained using the inverse of a known matrix, as shown in Eq. (26). The solution-dependent coefficient matrix $A_{d}$ is then calculated as shown in Eq. (27). Followed by one regular Jacobi iteration step with the new system matrix $A_{d}$ as shown in Eq. (28). Finally, the induced source $b_{i}$ is updated as shown in Eq. (29).
$x^{(k)}=A^{-1}\cdot(b+b_{i}^{(k)})$ (26)
$A_{d}^{(k)}=\digamma(x^{(k)})$ (27)
$x^{(k+1)}=-(A_{d}^{(k)}-I)\cdot x^{(k)}+b$ (28)
$b_{i}^{(k+2)}=\omega\cdot b_{i}^{(k)}+(1-\omega)\cdot(A-A_{d}^{(k)})\cdot x^{(k+1)}$ (29)
where $A^{-1}$ is a known inverse, calculated using the higher order method or other techniques, $A_{d}$ is the coefficient matrix dependent on the solution for the new system, $\digamma$ is a physics defined function that reflects the dependency of the coefficient matrix on the solution, and $\omega$ is a scalar relaxation factor that facilitates smoother convergence.
Eqs. (26) - (29) involve only matrix-vector multiplications, eliminating computationally expensive matrix-matrix multiplications. The equations are applied iteratively until convergence is achieved within a predefined tolerance, in a manner similar to a recurrent network structure where $\digamma$ plays the role of a nonlinear transformation.
3 Results
The time required for preparation, computation, and the overall process of solving matrix equations using built-in dense and sparse implementations of NumPy, SciPy, and PyTorch are presented in Fig. 1. Each data point is an average of 10 repetitions. The time consumed is calculated for the varying sizes of the matrices $n$ and are plotted on log-log scale. The last two data points of each curve are chosen for a regression fit on the logarithmic scale. The corresponding coefficients and exponents from the fit are mentioned on top of each figure for each implementation.
Similarly, the results for generating higher-order coefficients using matrix-matrix multiplication are presented in Fig. 2. Finally, the results for solving the matrix equation using the generated higher-order coefficient matrices using less computationally expensive matrix vector multiplication are shown in Fig. 3. The condensed results from Figs. 1-3 are presented in Table 1.
These results are computed using a system with Intel Core i7-10870H processor with 16 logic cores, 32 GB RAM, NVIDIA GeForce RTX 3070 Laptop GPU, Ubuntu 24.04.1 LTS operating system, along with software versions python=3.11, numpy=2.1, scipy=1.14, and pytorch-cuda=12.1.
Figure 1: Preparation, computation, and total times of solving system of linear equations using built-in matrix solve methods of NumPy, SciPy, and PyTorch.
Figure 2: Preparation, computation, and total times of generating higher order coefficient matrices using built-in matrix-matrix multiplication methods of NumPy, SciPy, and PyTorch.
Figure 3: Preparation, computation, and total times of solving matrix equation using generated higher order coefficient matrices with NumPy, SciPy, and PyTorch.
Table 1: Performance of solving matrix equations using built-in solve methods of Numpy, Scipy, PyTorch CPU/GPU and the higher-order method using respective built-in matrix multiplication methods.
Dense Numpy Scipy
PyTorch
CPU
PyTorch
GPU
Solving system of linear equations using built-in methods
   Preparation $5.41e^{-08}n^{1.86}$ $2.89e^{-08}n^{2.16}$ $6.31e^{-09}n^{1.98}$ $9.77e^{-06}n^{0.14}$
   Computation $2.77e^{-08}n^{1.93}$ $2.05e^{-10}n^{2.81}$ $5.31e^{-12}n^{2.84}$ $5.07e^{-10}n^{2.10}$
   Total $7.83e^{-08}n^{1.89}$ $3.60e^{-09}n^{2.52}$ $5.15e^{-10}n^{2.50}$ $5.15e^{-10}n^{2.10}$
Generating higher order coefficient matrices
   Preparation $1.71e^{-08}n^{2.02}$ $1.00e^{-04}n^{0.39}$ $1.26e^{-07}n^{1.46}$ $1.59e^{-05}n^{0.05}$
   Computation $3.38e^{-09}n^{3.31}$ $9.49e^{-05}n^{1.98}$ $1.05e^{-06}n^{2.45}$ $2.69e^{-05}n^{2.02}$
   Total $3.35e^{-09}n^{3.31}$ $9.67e^{-05}n^{1.97}$ $1.05e^{-06}n^{2.45}$ $2.69e^{-05}n^{2.02}$
Solving matrix equation using higher order coefficients
   Preparation $1.71e^{-08}n^{2.02}$ $1.00e^{-04}n^{0.39}$ $1.26e^{-07}n^{1.46}$ $1.59e^{-05}n^{0.05}$
   Computation $1.89e^{-08}n^{2.56}$ $6.90e^{-05}n^{1.46}$ $1.41e^{-05}n^{1.26}$ $1.06e^{-04}n^{1.20}$
   Total $2.51e^{-08}n^{2.42}$ $7.82e^{-05}n^{1.43}$ $1.40e^{-05}n^{1.26}$ $1.06e^{-04}n^{1.20}$
Sparse Numpy Scipy
PyTorch
CPU
PyTorch
GPU
Solving system of linear equations using built-in matrix solve methods
   Preparation $1.17e^{-06}n^{1.06}$ $4.33e^{-07}n^{1.17}$ $4.65e^{-07}n^{1.27}$ $1.70e^{-06}n^{1.04}$
   Computation $3.67e^{-10}n^{2.43}$ $4.59e^{-08}n^{1.42}$ $1.26e^{-11}n^{2.72}$ $8.52e^{-10}n^{2.03}$
   Total $4.97e^{-10}n^{2.40}$ $2.05e^{-08}n^{1.33}$ $6.47e^{-11}n^{2.54}$ $1.24e^{-08}n^{1.76}$
Generating higher order coefficient matrices
   Preparation $4.05e^{-08}n^{1.48}$ $1.08e^{-04}n^{0.18}$ $1.41e^{-05}n^{0.65}$ $1.51e^{-05}n^{0.63}$
   Computation $2.22e^{-09}n^{3.38}$ $9.92e^{-06}n^{2.68}$ $3.82e^{-06}n^{2.22}$ $6.77e^{-05}n^{1.86}$
   Total $2.22e^{-09}n^{3.38}$ $1.00e^{-05}n^{2.68}$ $3.83e^{-06}n^{2.22}$ $6.78e^{-05}n^{1.86}$
Solving matrix equation using higher order coefficients
   Preparation $4.05e^{-08}n^{1.48}$ $1.08e^{-04}n^{0.18}$ $1.41e^{-05}n^{0.65}$ $1.51e^{-05}n^{0.63}$
   Computation $3.36e^{-08}n^{2.33}$ $1.26e^{-05}n^{1.83}$ $2.17e^{-06}n^{1.57}$ $1.71e^{-04}n^{1.12}$
   Total $3.63e^{-08}n^{2.32}$ $1.42e^{-05}n^{1.80}$ $2.60e^{-06}n^{1.54}$ $1.74e^{-04}n^{1.12}$
4 Discussions
The computational complexity of solving matrix equations depends primarily on two factors [35, 36, 37], the sparsity of the coefficient matrices and the type of method used, direct or iterative. The system matrices built using PDEs are usually sparse with very few nonzero elements in each row and are diagonally dominant. The computational complexity of solving such sparse systems is much lower when compared with their denser counterparts, such as Hamiltonian matrices in computational quantum chemistry, due to the reduced number of calculations that can be performed in each row. A comprehensive comparison of different computational complexities is shown in Table 2.
Table 2: Comparison of computational complexities of solving matrix equations and matrix multiplications using different known methods.
Method Complexity
Solving Matrix Equation
   Gaussian Elimination (Dense) $O(n^{3})$
   LU Decomposition (Dense) $O(n^{3})$
   Jacobi Iteration (Dense, $k\sim n$) $O(n^{3})$
   Jacobi Iteration (Sparse, $s\ll n$, $k\ll n$) [38] $O(kns)$
   Conjugate Gradient (SPD matrix) [39, 22] $O(n^{2})$ to $O(n\log n)$
   Multigrid Method [40] $O(n)$
Matrix-Matrix Multiplication
   Naive Matrix Multiplication $O(n^{3})$
   Strassen’s Algorithm [41] $O(n^{\log_{2}(7)})\approx O(n^{2.81})$
   Coppersmith-Winograd Algorithm [42] $O(n^{2.376})$
   Improved Variants of Coppersmith-Winograd [43] $O(n^{2.3728596})$
Matrix-Vector Multiplication $O(n^{2})$
The proposed higher-order method has the same computational complexity as matrix-matrix multiplication. Even though it is an order of magnitude higher than the matrix-vector multiplication, unlike iterative operations, the matrix-matrix multiplication is a highly parallelizable operation [44, 45, 46, 47].
Moreover, the theoretical computational complexity of multiplying large matrices has improved significantly over the years, falling from default $O(n^{3})$ to the current known limit of $O(n^{2.376})$ [42] for large matrices.
From Figs. 1, 2, and 3, it is clear that the actual computation times of the higher-order method while generating the higher-order coefficients and even solving matrix equations using the generated coefficients are higher than the built-in matrix equation solving methods of the software packages used. However, the order of complexity of generating coefficients is better than the built-in matrix solve methods of any package, which suggests an algorithmic improvement in solving matrix equations. This also suggests that the proposed higher-order method is highly parallelized, making it exceptionally well suited for leveraging advancements in parallel computing.
As coefficient matrices become denser with recursive squaring, pruning techniques [48, 49] used in traditional neural networks can be adopted to make dense higher-order matrices much more compact. However, with recent developments in parallel computing and the ever-increasing affordability of GPUs [50, 51] and even photonic circuits [52], matrix-matrix multiplications can be accelerated.
The proposed higher-order method for matrix inversion structurally aligns with a deep linear network. Once the coefficients are explicitly derived for a known system, the method enables efficient resolution of variations of that system without the need for computationally intensive re-computation of the coefficients. The recurrent iterative scheme inherently adopts the form of a deep recurrent neural network. This framework aims to provide analytical insight into the explicit definition of network parameters, thereby circumventing the training process and significantly reducing computational overhead.
5 Conclusion
Physical systems are fundamentally governed by differential equations, and recent advances have shown growing interest in Physics-Informed Neural Networks (PINNs) for solving such equations. However, in most existing frameworks, the term “physics-informed” pertains primarily to the incorporation of physical laws into the training objective, rather than to the explicit definition of network parameters such as weights, biases, and activation functions.
This work presents a direction toward interpreting and defining the inner workings of neural networks through explicit analytical formulation. In particular, a modified Jacobi iteration method is developed to compute higher-order coefficient matrices using matrix-matrix multiplications. The resulting formulation, referred to as the higher-order Jacobi method, exhibits the structural characteristics of a shallow linear network with a limited number of layers. Through this formulation, it is also possible to compute the inverse of the original coefficient matrix directly via matrix operations, eliminating the need for iterative inversion procedures.
In addition, an iterative solution scheme is proposed whereby the network parameters, once determined for a known system, can be reused to solve variations of the system without re-computing the coefficients. This capability is especially advantageous in applications where the underlying system undergoes parametric changes. The iterative process takes the form of a deep recurrent neural network, offering a principled and efficient approach that maintains consistency across related systems.
Computational analysis indicates that the proposed methods lead to a measurable improvement in the order of complexity. This signifies a potential algorithmic advancement in solving matrix equations more efficiently, with reduced computational cost and improved scalability. The presented methodology not only connects traditional numerical methods with neural architectures but also contributes toward interpretable [53, 54] and training-free network designs grounded in physical and mathematical principles.
Future directions include the extension of this framework to non-linear systems and time-dependent differential equations, enabling broader applications in scientific computing and model-driven learning.

### v2

@@
-Higher order Jacobi method for solving a system of linear equations
+Higher order Jacobi method for solving system of linear equations
 Abstract
 This work proposes a higher-order iterative framework for solving matrix equations, inspired by the structure and functionality of neural networks.
-A modification of the classical Jacobi method is introduced to compute higher-order coefficient matrices through matrix-matrix multiplications.
-The resulting method, termed the higher-order Jacobi method, structurally resembles a shallow linear network and allows direct computation of the inverse of the coefficient matrix.
-Building on this, an iterative scheme is developed that, once the network parameters are determined for a known system, enables efficient resolution of system variations without re-computing the coefficients.
+A modification of the classical Jacobi iterative method is introduced to compute higher-order coefficient matrices through matrix-matrix multiplications.
+The resulting method, termed the higher order Jacobi method (HOJM), structurally resembles a shallow linear network and allows direct computation of the inverse of the coefficient matrix.
+Building on this, an iterative scheme is developed that allows efficient resolution of system variations without recomputing the coefficients, once the network parameters are trained for a known system.
 This iterative process naturally assumes the form of a deep recurrent neural network.
-The proposed approach moves beyond conventional Physics-Informed Neural Networks (PINNs) by providing an explicit, training-free definition of network parameters rooted in physical and mathematical formulations.
-Computational analysis demonstrates improved order of complexity, suggesting a promising direction for algorithmic efficiency in solving linear systems.
+The proposed approach goes beyond conventional physics-informed neural networks (PINNs) by providing an explicit, training-free definition of network parameters rooted in physical and mathematical formulations.
+Computational analysis on GPU reveals significant enhancement in the order of complexity, highlighting a compelling and transformative direction for advancing algorithmic efficiency in solving linear systems.
 This methodology opens avenues for interpretable and scalable solutions to physically motivated problems in computational science.
 1 Introduction
-Many real-world phenomena, ranging from fluid dynamics and heat transfer to financial modeling and electromagnetism, are governed by Partial Differential Equations (PDEs).
+Many real-world phenomena, ranging from fluid dynamics and heat transfer to financial modeling, electromagnetism, and more, are governed by Partial Differential Equations (PDEs).
 These equations describe how physical quantities evolve over space and time.
@@
 To enable numerical solutions, partial differential equations (PDEs) [1, 2, 3] are discretized, converting the continuous problem into a system of algebraic equations through finite-difference or finite-element approximations.
-This discretization process leads to a matrix equation of the form shown in Eq.
+This discretization process often leads to a simple matrix equation of the form shown in Eq.
 (1).
@@
 where $A$ is a system matrix derived from discretization, $x$ is the vector of unknowns, and $b$ represents known constraints or forcing terms.
-Efficiently solving this system is fundamental in computational science [4, 5], enabling simulations [6, 7, 8] and optimizations [9, 10] across various engineering and scientific domains.
+Efficiently solving this system is fundamental in computational science [4, 5], enabling simulations [6, 7, 8] and optimizations [9, 10] in various engineering and scientific domains.
+Although the discretization of PDEs can also lead to eigenvalue equations, this article will focus exclusively on a simpler matrix equation.
 There are several approaches to solving the matrix equation.
 Direct methods [11, 12, 13], such as Gaussian elimination [14] and LU decomposition [15], solve for $x$ exactly when feasible.
-Iterative methods [16, 17], such as the Jacobi [18, 19] and conjugate gradient methods [20, 21, 22], are effective for large, sparse systems.
-The AI driven approach, physics-informed neural networks (PINNs) [23, 24, 25], leverages deep learning to approximate solutions while respecting underlying physical laws.
+Iterative methods [16, 17], such as Jacobi [18, 19] and conjugate gradient methods [20, 21, 22], are effective for large sparse systems.
+The AI driven approach, physics-informed neural networks (PINNs) [23, 24, 25], leverages deep learning to approximate solutions while respecting the underlying physical laws.
 This article presents a variation of the Jacobi method that can harness emerging hardware, such as GPUs being developed for AI to solve matrix equations.
@@
 Notably, this higher-order method (HOM) can also be seamlessly extended to handle dense matrices, a scenario where traditional Jacobi methods are typically less effective.
-Although the current state of physics-informed neural networks largely emphasizes physics-informed training [26, 27], our work shifts focus toward intuitive, physics-informed explicit generation of weights and biases, paving the way for a more interpretable framework where the physics itself defines the structure of the network.
+Although the current state of physics-informed neural networks largely emphasizes physics-informed training [26, 27], our work shifts focus to intuitive, physics-informed explicit generation of weights and biases, paving the way for a more interpretable framework where the physics itself defines the structure of the network.
 This study demonstrates explicit weight and bias construction for linear systems.
@@
 2 The Higher Order Method
-The Jacobi approach to solving a matrix equation follows a decomposition of the matrix $A$ into its diagonal $D$, lower $L$, and upper $U$ matrices.
-The solution is then iteratively computed until the solution converges within a desirable tolerance, as shown in Eq.
-(2).
+The Jacobi approach to solving a matrix equation follows a decomposition of the matrix $A$ into its diagonal $D$, lower $L$ and upper $U$ matrices.
+The solution is then iteratively computed as shown in Eq.
+(2), until the solution converges within a desirable tolerance.
 $x^{(k+1)}=D^{-1}\cdot(b-(L+U)\cdot x^{(k)})$ (2)
-where $k$ and $k+1$ represent the current and next iteration steps in the iterative process.
+where $k$ and $k+1$ represent the current and next solution guesses respectively in the iterative process for $k\geq 0$.
 2.1 The higher order Jacobi method for solving matrix equation
@@
 where $I$ is the identity matrix. $A_{n}$ and $b_{n}$ are normalized matrices of $A$ and $b$, respectively.
-Further simplification of Eq. (3) to express the next iteration as a linear combination of the current one leads to Eq. (4).
+Further simplification of Eq. (3) to express the next guess $x^{(k+1)}$ as a linear combination of the current guess $x^{(k)}$ leads to Eq. (4).
 $x^{(k+1)}=\alpha\cdot x^{(k)}+\beta$ (4)
@@
 $\beta=b_{n}$ (6)
-This transformation ensures a more structured and scalable computation.
-Starting with an initial guess $x^{(0)}$, the solution after the first iteration $x^{(1)}$ is calculated as shown in Eq. (7)
+This transformation ensures structured and scalable computation of the higher order coefficient matrices.
+Starting with an initial guess $x^{(0)}$, the solution after the first iteration $x^{(1)}$ is calculated as shown in Eq. (7).
 $x^{(1)}=\alpha\cdot x^{(0)}+\beta$ (7)
-Similarly, the second iteration is computed as shown in Eq. (8)
+Similarly, the solution in second iteration $x^{(2)}$ is computed using the solution from the first iteration $x^{(1)}$ as shown in Eq. (8).
 $x^{(2)}=\alpha\cdot x^{(1)}+\beta$ (8)
-However, substituting Eq. (7) into Eq. (8), corresponding to the second-order approximation, enables direct computation of the second iterate from the initial guess, resulting in the following.
+However, substituting Eq.
+(7) into Eq.
+(8), corresponding to the second-order approximation, enables direct computation of the second iterate $x^{(2)}$ from the initial guess $x^{(0)}$, resulting in the following:
 $x^{(2)}=\alpha\cdot(\alpha\cdot x^{(0)}+\beta)+\beta$ (9)
@@
 $x^{(2)}=\alpha^{2}\cdot x^{(0)}+(\alpha+I)\cdot\beta$ (10)
-The simplification of Eq. (10) yields Eq. (11), explicitly illustrating the linear dependency of the second iterate on the initial guess.
+Simplification of Eq. (10) yields Eq. (11), explicitly illustrating the linear dependency of the second iterate $x^{(2)}$ on the initial guess $x^{(0)}$.
 $x^{(2)}=\alpha^{(1)}\cdot x^{(0)}+\beta^{(1)}$ (11)
@@
 $\beta^{(1)}=(\alpha+I)\cdot\beta$ (13)
-Further generalizing Eq. (11),
-$x^{(2k)}=\alpha^{(1)}\cdot x^{(0)}+\beta^{(1)}$ (14)
-which allows iterations to be performed using the second-order coefficients, $\alpha^{(1)}$ and $\beta^{(1)}$, respectively, instead of the first-order coefficients $\alpha$ and $\beta$.
+Further generalizing Eq. (11), results in Eq. (14)
+$x^{(2k+2)}=\alpha^{(1)}\cdot x^{(2k)}+\beta^{(1)}$ (14)
+which allows iterations to be performed using the second-order coefficients, $\alpha^{(1)}$ and $\beta^{(1)}$, respectively, instead of the first-order coefficients $\alpha$ and $\beta$ for $k\geq 0$.
 Consequently, a single iteration with the second-order coefficients is equivalent to two iterations with the first-order coefficients.
@@
 (11) takes the same form as Eq.
-(7), i.e., having the linear dependence on the initial guess.
-Following the same procedure of calculating second higher order coefficients, recursively substituting the higher order equation into itself, the general form of the higher order Jacobi method is represented in Eq.
+(7), that is, it has a linear dependence on the initial guess.
+Following the same procedure of calculating second higher-order coefficients, recursively substituting the higher order equation into itself, the general form of the higher order Jacobi method (HOJM) is presented in Eq.
 (15).
@@
 Consequently, the solution after $2^{k}$ iterations with first-order coefficients is equivalent to 1 iteration with the $k^{th}$-order coefficients.
-This exponential reduction in the number of iterations comes at the cost of increased computational complexity.
+Using Eq.
+(15), any higher order guess $x^{(2^{k})}$ can be calculated directly from the initial guess $x^{(0)}$.
+However, each higher-order guess $x^{(2^{k})}$ can also be obtained from the previous higher-order guess $x^{(2^{k-1})}$ as shown in Eq.
+(18), for faster convergence.
+$x^{(2^{k})}=\alpha^{(k)}\cdot x^{(2^{k-1})}+\beta^{(k)}$ (18)
+The exponential reduction in the number of iterations comes at the cost of an increase in the computational complexity of matrix-matrix multiplication.
 Each higher-order step requires multiplying two square matrices, whereas traditional Jacobi iterations only involve multiplying a square matrix with a column vector.
 For large matrices, block matrix multiplication and GPU acceleration with parallel processing significantly reduce real-world execution time of computing the higher-order coefficients.
-2.2 Finding inverse of coefficient matrix from the higher order Jacobi method
+2.2 Finding inverse of coefficient matrix using the higher order Jacobi method
 The expanded form of Eq.
@@
 It is well-established that any linear network, irrespective of its depth, can be equivalently represented as a single-layer network.
-In the context of the proposed higher-order Jacobi method, this single-layer equivalent corresponds to the inverse of the normalized coefficient matrix $A_{n}$, as elaborated in the following discussion.
+In the context of the proposed higher order Jacobi method, this single-layer equivalent corresponds to the inverse of the normalized coefficient matrix $A_{n}$, as elaborated in the following discussion.
 In Eq.
-(15), the solution uniquely depends on the bias term $\beta^{(k-1)}$, while the other term is arbitrary, even though $\alpha^{(k-1)}$ is uniquely defined, as the initial guess can be arbitrary.
-The term $\alpha$ times $x$ can become insignificant with null vector as an initial guess.
+(15), the solution uniquely depends on the bias term $\beta^{(k)}$, while the other term is arbitrary, as the initial guess can be arbitrary, even though $\alpha^{(k)}$ is uniquely defined.
+The term $\alpha$ times $x$ can become insignificant with a null vector as an initial guess.
 Thus, the multiplication term can be omitted, resulting in Eq.
@@
 The explicit closed-form representation of the inverse of matrix $A_{n}$, as shown in Eq.
-(24), is derived from an iterative Jacobi method, which proves the equivalence of direct and iterative methods of solving the matrix equation, whenever necessary conditions for convergence are met in the regular Jacobi iteration method [28, 29, 30].
-In addition, the order of accuracy can be selected; the higher the order of coefficient matrices, the higher the accuracy.
+(25), is derived from the Jacobi method, which proves the equivalence of direct and iterative methods of solving the matrix equation, when necessary conditions for convergence [28, 29, 30] are met in the regular Jacobi iteration method.
+In addition, the order of accuracy can also be selected; higher the order of coefficient matrices, higher the accuracy.
 Eq.
@@
 Although Eq. (18) takes the same form as a deep linear network with weights and biases, Eq. (22) represents its transformation into a single-layer linear network, especially one without biases.
-For an arbitrary forcing matrix $b$, the higher-order method produces a solution similar to the regular Jacobi iteration method.
-However, in real-world applications, the system matrix $A$ can be altered due to the material properties embedded in different PDEs [31] or due to mesh refinement [32], which requires the recalculation of higher-order matrices, which is computationally expensive.
-This recalculation, even if there is a single change in one value of the original coefficient matrix, has the same computational complexity as matrix-matrix multiplication, making it inefficient for dynamically changing systems, especially for non-linear systems, where coefficients depend on the solution.
-This is where an iterative approach finds its way in and gives insight into importance/need of layers in neural networks.
-One such process is discussed in the following section.
+For an arbitrary forcing matrix $b$, the higher-order method produces a solution the same as the regular Jacobi iteration method in the equivalent number of iterations.
+However, in real-world applications, the system matrix $A$ is not constant and can change from problem to problem due to the material properties embedded in different PDEs [31] or due to mesh refinement [32], which requires recalculation of the higher-order matrices, which may be computationally intensive.
+Even with single change in a value of the original coefficient matrix, involve recalculation of the higher-order coefficient matrices using matrix-matrix multiplication, making it inefficient for dynamically changing systems, especially for non-linear systems, where values of the coefficient matrix depend on the solution.
+At this stage, iterative methods constitute the only viable pathway for rigorously understanding the structural necessity and functional dynamics of layered architectures in neural networks.
+A representative process is presented in detail in the following section.
 2.4 Induced source iterative scheme as an explicit recurrent neural network
-While the coefficient matrix may change dramatically in some of the situations such as mesh refinement, the following scheme is proposed for slowly varying coefficient matrix such as in nonlinear systems and even eigenvalue problems.
-Instead of recalculating the coefficients, the goal is to modify the source matrix such that solving the modified source matrix [33, 34] with a known inverse produces the same result as solving the desired system without recalculating its inverse.
+The coefficient matrix $A$ may undergo significant changes in scenarios such as mesh refinement.
+However, the following scheme is designed for cases where the matrix varies gradually, as encountered in nonlinear systems.
+Instead of recalculating the modified higher-order coefficients, the goal is to adjust the source matrix [33, 34] so that solving a system with unmodified higher-order coefficients and modified source yields the same solution as solving the system with updated coefficients and the original source.
 This approach eliminates the need for recomputing higher-order coefficients while maintaining solution consistency.
-Let the modified source be $b_{m}$ as shown in Eq. (25).
+Let the modified source $b_{m}$ consist of the original source $b$ and the induced source $b_{i}$ as defined in Eq. (26).
 $b_{m}=b+b_{i}$ (26)
-where $b$ is the original source matrix for which the new system will be solved and $b_{i}$ is the induced source that will be computed iteratively.
-This iterative approach introduces the following series of intermediate steps in which the solution is first obtained using the inverse of a known matrix, as shown in Eq.
+where $b$ is the original source matrix for which the new system is to be solved and $b_{i}$ is the induced source, initialized as a null matrix and computed iteratively.
+This iterative approach introduces the following series of intermediate steps in which the solution is first obtained using the known inverse of the coefficient matrix, as shown in Eq.
 (27).
@@
 $b_{i}^{(k+2)}=\omega\cdot b_{i}^{(k)}+(1-\omega)\cdot(A-A_{d}^{(k)})\cdot x^{(k+1)}$ (30)
-where $A^{-1}$ is a known inverse, calculated using the higher order method or other techniques, $A_{d}$ is the coefficient matrix dependent on the solution for the new system, $\digamma$ is a physics defined function that reflects the dependency of the coefficient matrix on the solution, and $\omega$ is a scalar relaxation factor that facilitates smoother convergence.
+where $A^{-1}$ is a known inverse, calculated using the higher order method or other techniques, $A_{d}$ is the coefficient matrix dependent on the solution for the new system, $\digamma$ is a physics defined function that reflects the dependency of the coefficient matrix on the solution, that is, the nonlinearity, and $\omega$ is a scalar relaxation factor that facilitates smoother convergence.
 Eqs.
 (27) - (30) involve only matrix-vector multiplications, eliminating computationally expensive matrix-matrix multiplications.
-The equations are applied iteratively until convergence is achieved within a predefined tolerance, in a manner similar to a recurrent network structure where $\digamma$ plays the role of a nonlinear transformation.
+These equations are applied iteratively until convergence is achieved with a predefined tolerance, similar to a recurrent network structure where $\digamma$ plays the role of the nonlinear transformation.
 3 Results
-The time required for preparation, computation, and the overall process of solving matrix equations using built-in dense and sparse implementations of NumPy, SciPy, and PyTorch are presented in Fig.
+A comparison of the time taken to solve a matrix equation and to invert a matrix with increasing matrix size using different methods is shown in Fig.
 1.
-Each data point is an average of 10 repetitions.
-The time consumed is calculated for the varying sizes of the matrices $n$ and are plotted on log-log scale.
-The last two data points of each curve are chosen for a regression fit on the logarithmic scale.
-The corresponding coefficients and exponents from the fit are mentioned on top of each figure for each implementation.
-Similarly, the results for generating higher-order coefficients using matrix-matrix multiplication are presented in Fig.
-2.
-Finally, the results for solving the matrix equation using the generated higher-order coefficient matrices using less computationally expensive matrix vector multiplication are shown in Fig.
-3.
-The condensed results from Figs.
-1-3 are presented in Table 1.
-These results are computed using a system with Intel Core i7-10870H processor with 16 logic cores, 32 GB RAM, NVIDIA GeForce RTX 3070 Laptop GPU, Ubuntu 24.04.1 LTS operating system, along with software versions python=3.11, numpy=2.1, scipy=1.14, and pytorch-cuda=12.1.
-Figure 1: Preparation, computation, and total times of solving system of linear equations using built-in matrix solve methods of NumPy, SciPy, and PyTorch.
-Figure 2: Preparation, computation, and total times of generating higher order coefficient matrices using built-in matrix-matrix multiplication methods of NumPy, SciPy, and PyTorch.
-Figure 3: Preparation, computation, and total times of solving matrix equation using generated higher order coefficient matrices with NumPy, SciPy, and PyTorch.
-Table 1: Performance of solving matrix equations using built-in solve methods of Numpy, Scipy, PyTorch CPU/GPU and the higher-order method using respective built-in matrix multiplication methods.
-Dense Numpy Scipy
+The results in Fig.
+1(a) & 1(c) were obtained on a computer system (System 1) with lower specifications, while those in Fig.
+1(b) & 1(d) were obtained on a higher-end system (System 2).
+The configurations of both systems are presented in Table 1.
+The matrix equation is solved using the built-in solve method from the linear algebra subpackage of each Python library for the existing methods, and using the built-in matrix-matrix multiplication functionality in PyTorch for the higher order Jacobi method implementation.
+The matrix inverse is similarly computed using the respective built-in inverse functions for existing methods, while the higher order Jacobi method approximates the inverse through iterative matrix-matrix operations.
+The reported inverse times also account for the re-normalization of the coefficient product.
+The x-axis represents the matrix size, while the y-axis shows the time taken to compute the solution.
+The results are presented in log-log scale.
+All the calculations are performed with 32-bit single presision data formats.
+The horizontal extent of each curve depends on the available system memory for the corresponding operation, whereas the vertical range is capped to 20 seconds for any single operation whether it is solving a single matrix equation or inverting a matrix.
+The coeffecient matrix $A$ is constructed to be sparse.
+Irrespective of the size of the matrix, 20 higher coeffecients are calculated while producing the higher order Jacobi results.
+20 higher order iterations is equivalent to 1 million regular Jacobi iterations.
+The accuracy in the solution from 20 higher order iterations is equivalent to accuracy of Jacobi method after million iterations.
+The CPU implementations of the methods are shown using solid lines, while the GPU implementations are represented by dotted lines.
+In System 1, a Conda environment was created, and the required Python packages, as listed in Table 1, were installed.
+In contrast, System 2 utilizes a default docker pytorch container, as it is part of a high-performance GPU cluster.
+It is due to this reason the package versions are slightly different and scipy is not available on System 2.
+Since System 2 is primarily designed for GPU computing, a dedicated GPU is allocated when the code is executed, whereas its CPU may have been reserved for other tasks.
+The runtime empirical computational complexity for each curve is calculated using the last two data points.
+If a sudden jump is observed in the last 3 data points, the middle point is excluded for the calculation.
+The corresponding results are presented in Table 2.
+The best order in each row has be highlighted.
+From Fig.
+1 and Table 2, it is evident that the higher order Jacobi implementation using PyTorch GPU tensors performs the best, regardless of the system or whether a solve or inverse method is used.
+This performance advantage is attributed to the higher order Jacobi method’s ability to leverage the parallelism inherent in matrix-matrix multiplications.
+Figure 1: Performance Comparision of built-in methods and the proposed higher order Jacobi method in solving system of linear equations and finding matrix inverse, (a) Solving matrix equation on System 1, (b) Solving matrix equation on System 2, (c) Finding matrix inverse on System 1, and (d) Finding matrix inverse on System 2.
+Table 1: Configuration of the two compute systems used to compare the performance.
+System 1 System 2
+CPU Intel Core i7-10870H Intel Xeon Sapphire Rapids
+Cores 8 2 $\times$ 56
+System RAM 32 GB 2 TB
+GPU
+NVIDIA GeForce RTX 3070
+Laptop Version
+NVIDIA Hopper H100
+Tensor Core
+Graphics Memory 8 GB 80 GB
+Operating System Ubuntu 24.04.1 LTS
+Red Hat Enterprise
+Linux (RHEL) 9.2
+Python Package Versions
+python = 3.10.18, numpy = 2.2.6,
+pytorch = 2.5.1, pytorch-cuda = 11.8,
+scipy = 1.15.2
+python = 3.10.13, numpy = 1.26.3,
+pytorch = 2.2.1, pytorch-cuda = 12.1
+Table 2: Performance of solving matrix equation and finding inverse using built-in methods of Numpy, Scipy, PyTorch CPU/GPU and the higher order Jacobi method using PyTorch built-in matrix-matrix multiplication.
+NumPy
+linalg
+Scipy
+linalg
 Pytorch
-CPU
+linalg
 Pytorch
-GPU
-Solving system of linear equations using built-in methods
-   Preparation $5.41e^{-08}n^{1.86}$ $2.89e^{-08}n^{2.16}$ $6.31e^{-09}n^{1.98}$ $9.77e^{-06}n^{0.14}$
-   Computation $2.77e^{-08}n^{1.93}$ $2.05e^{-10}n^{2.81}$ $5.31e^{-12}n^{2.84}$ $5.07e^{-10}n^{2.10}$
-   Total $7.83e^{-08}n^{1.89}$ $3.60e^{-09}n^{2.52}$ $5.15e^{-10}n^{2.50}$ $5.15e^{-10}n^{2.10}$
-Generating higher order coefficient matrices
-   Preparation $1.71e^{-08}n^{2.02}$ $1.00e^{-04}n^{0.39}$ $1.26e^{-07}n^{1.46}$ $1.59e^{-05}n^{0.05}$
-   Computation $3.38e^{-09}n^{3.31}$ $9.49e^{-05}n^{1.98}$ $1.05e^{-06}n^{2.45}$ $2.69e^{-05}n^{2.02}$
-   Total $3.35e^{-09}n^{3.31}$ $9.67e^{-05}n^{1.97}$ $1.05e^{-06}n^{2.45}$ $2.69e^{-05}n^{2.02}$
-Solving matrix equation using higher order coefficients
-   Preparation $1.71e^{-08}n^{2.02}$ $1.00e^{-04}n^{0.39}$ $1.26e^{-07}n^{1.46}$ $1.59e^{-05}n^{0.05}$
-   Computation $1.89e^{-08}n^{2.56}$ $6.90e^{-05}n^{1.46}$ $1.41e^{-05}n^{1.26}$ $1.06e^{-04}n^{1.20}$
-   Total $2.51e^{-08}n^{2.42}$ $7.82e^{-05}n^{1.43}$ $1.40e^{-05}n^{1.26}$ $1.06e^{-04}n^{1.20}$
-Sparse Numpy Scipy
-PyTorch
-CPU
-PyTorch
-GPU
-Solving system of linear equations using built-in matrix solve methods
-   Preparation $1.17e^{-06}n^{1.06}$ $4.33e^{-07}n^{1.17}$ $4.65e^{-07}n^{1.27}$ $1.70e^{-06}n^{1.04}$
-   Computation $3.67e^{-10}n^{2.43}$ $4.59e^{-08}n^{1.42}$ $1.26e^{-11}n^{2.72}$ $8.52e^{-10}n^{2.03}$
-   Total $4.97e^{-10}n^{2.40}$ $2.05e^{-08}n^{1.33}$ $6.47e^{-11}n^{2.54}$ $1.24e^{-08}n^{1.76}$
-Generating higher order coefficient matrices
-   Preparation $4.05e^{-08}n^{1.48}$ $1.08e^{-04}n^{0.18}$ $1.41e^{-05}n^{0.65}$ $1.51e^{-05}n^{0.63}$
-   Computation $2.22e^{-09}n^{3.38}$ $9.92e^{-06}n^{2.68}$ $3.82e^{-06}n^{2.22}$ $6.77e^{-05}n^{1.86}$
-   Total $2.22e^{-09}n^{3.38}$ $1.00e^{-05}n^{2.68}$ $3.83e^{-06}n^{2.22}$ $6.78e^{-05}n^{1.86}$
-Solving matrix equation using higher order coefficients
-   Preparation $4.05e^{-08}n^{1.48}$ $1.08e^{-04}n^{0.18}$ $1.41e^{-05}n^{0.65}$ $1.51e^{-05}n^{0.63}$
-   Computation $3.36e^{-08}n^{2.33}$ $1.26e^{-05}n^{1.83}$ $2.17e^{-06}n^{1.57}$ $1.71e^{-04}n^{1.12}$
-   Total $3.63e^{-08}n^{2.32}$ $1.42e^{-05}n^{1.80}$ $2.60e^{-06}n^{1.54}$ $1.74e^{-04}n^{1.12}$
+HOJM
+Solving Matrix
+Equation
+CPU GPU CPU GPU
+while
+System 1 $9.5e^{-12}n^{2.86}$ $3.4e^{-08}n^{1.28}$ $2.5e^{-13}n^{3.15}$ $1.2e^{-08}n^{1.80}$ $2.0e^{-11}n^{3.32}$ $2.4e^{-04}n^{0.19}$
+System 2 $7.2e^{-11}n^{2.78}$ $3.9e^{-12}n^{3.00}$ $1.8e^{-13}n^{2.78}$ $5.6e^{-10}n^{3.04}$ $4.8e^{-06}n^{0.79}$
+after
+System 1 $2.2e^{-08}n^{1.98}$ $6.1e^{-04}n^{0.01}$
+System 2 $7.5e^{-12}n^{3.02}$ $1.4e^{-05}n^{0.37}$
+Finding Matrix
+Inverse
+CPU GPU CPU GPU
+System 1 $2.6e^{-11}n^{2.93}$ $4.7e^{-07}n^{1.81}$ $3.0e^{-13}n^{3.33}$ $1.6e^{-11}n^{2.61}$ $2.0e^{-09}n^{2.88}$ $2.5e^{-04}n^{0.24}$
+System 2 $1.3e^{-10}n^{2.85}$ $3.1e^{-12}n^{3.18}$ $1.3e^{-12}n^{2.71}$ $3.2e^{-09}n^{2.93}$ $1.8e^{-03}n^{-0.07}$
 4 Discussions
@@
 The system matrices built using PDEs are usually sparse with very few nonzero elements in each row and are diagonally dominant.
-The computational complexity of solving such sparse systems is much lower when compared with their denser counterparts, such as Hamiltonian matrices in computational quantum chemistry, due to the reduced number of calculations that can be performed in each row.
+The computational complexity of solving such sparse systems is much lower when compared with their denser counterparts, such as Hamiltonian matrices in computational quantum chemistry.
+This is due to the reduced number of calculations performed in each row.
 A comprehensive comparison of different computational complexities is shown in Table 3.
@@
    Strassen’s Algorithm [41] $O(n^{\log_{2}(7)})\approx O(n^{2.81})$
+   Google’s AlphaEvolve [42] $O(n^{\log_{4}(48)})\approx O(n^{2.79})$
    Coppersmith-Winograd Algorithm [43] $O(n^{2.376})$
@@
 The proposed higher-order method has the same computational complexity as matrix-matrix multiplication.
-Even though it is an order of magnitude higher than the matrix-vector multiplication, unlike iterative operations, the matrix-matrix multiplication is a highly parallelizable operation [44, 45, 46, 47].
-Moreover, the theoretical computational complexity of multiplying large matrices has improved significantly over the years, falling from default $O(n^{3})$ to the current known limit of $O(n^{2.376})$ [42] for large matrices.
-From Figs.
-1, 2, and 3, it is clear that the actual computation times of the higher-order method while generating the higher-order coefficients and even solving matrix equations using the generated coefficients are higher than the built-in matrix equation solving methods of the software packages used.
-However, the order of complexity of generating coefficients is better than the built-in matrix solve methods of any package, which suggests an algorithmic improvement in solving matrix equations.
-This also suggests that the proposed higher-order method is highly parallelized, making it exceptionally well suited for leveraging advancements in parallel computing.
+Although its naive approach is an order of magnitude higher than matrix-vector multiplication, matrix-matrix multiplication is a highly parallelizable operation [45, 46, 47, 48].
+Moreover, the theoretical computational complexity of multiplying large matrices has improved significantly over the years, falling from the default $O(n^{3})$ to the current known limit of $O(n^{2.376})$ [43] for large matrices.
+However, practical and scalable implementations even with the help of deep learning algorithms [42, 49]still tend to perform more closely to the Strassen algorithm in practice.
+However, the higher order Jacobi method takes advantage of the parallelizability of matrix-matrix multiplication, unlike the current sequential approaches, and therefore demonstrates exceptional performance when implemented on GPUs.
 As coefficient matrices become denser with recursive squaring, pruning techniques [50, 51] used in traditional neural networks can be adopted to make dense higher-order matrices much more compact.
-However, with recent developments in parallel computing and the ever-increasing affordability of GPUs [50, 51] and even photonic circuits [52], matrix-matrix multiplications can be accelerated.
+However, with recent developments in parallel computing and the ever-increasing affordability of GPUs [52, 53] and even photonic circuits [54], matrix-matrix multiplications can be further accelerated.
 The proposed higher-order method for matrix inversion structurally aligns with a deep linear network.
@@
 The recurrent iterative scheme inherently adopts the form of a deep recurrent neural network.
-This framework aims to provide analytical insight into the explicit definition of network parameters, thereby circumventing the training process and significantly reducing computational overhead.
+This framework aims to provide analytical insight into the explicit definition of network parameters in physics informed neural networks, thereby circumventing the training process and significantly reducing computational overhead.
 5 Conclusion
 Physical systems are fundamentally governed by differential equations, and recent advances have shown growing interest in Physics-Informed Neural Networks (PINNs) for solving such equations.
-However, in most existing frameworks, the term “physics-informed” pertains primarily to the incorporation of physical laws into the training objective, rather than to the explicit definition of network parameters such as weights, biases, and activation functions.
-This work presents a direction toward interpreting and defining the inner workings of neural networks through explicit analytical formulation.
+However, in most existing frameworks, the term “physics-informed” relates primarily to the incorporation of physical laws into the training objective, rather than to the explicit definition of network parameters such as weights, biases, and activation functions.
+This work presents a direction toward interpreting and defining the inner workings of neural networks through an explicit analytical formulation.
 In particular, a modified Jacobi iteration method is developed to compute higher-order coefficient matrices using matrix-matrix multiplications.
-The resulting formulation, referred to as the higher-order Jacobi method, exhibits the structural characteristics of a shallow linear network with a limited number of layers.
+The resulting formulation, referred to as the higher order Jacobi method, exhibits the structural characteristics of a shallow linear network with a limited number of layers.
 Through this formulation, it is also possible to compute the inverse of the original coefficient matrix directly via matrix operations, eliminating the need for iterative inversion procedures.
-In addition, an iterative solution scheme is proposed whereby the network parameters, once determined for a known system, can be reused to solve variations of the system without re-computing the coefficients.
+In addition, an iterative scheme is proposed in which the network parameters, once determined for a known system, can be reused to solve variations of the system without recomputing the coefficients.
 This capability is especially advantageous in applications where the underlying system undergoes parametric changes.
 The iterative process takes the form of a deep recurrent neural network, offering a principled and efficient approach that maintains consistency across related systems.
-Computational analysis indicates that the proposed methods lead to a measurable improvement in the order of complexity.
-This signifies a potential algorithmic advancement in solving matrix equations more efficiently, with reduced computational cost and improved scalability.
-The presented methodology not only connects traditional numerical methods with neural architectures but also contributes toward interpretable [53, 54] and training-free network designs grounded in physical and mathematical principles.
-Future directions include the extension of this framework to non-linear systems and time-dependent differential equations, enabling broader applications in scientific computing and model-driven learning.
+Computational analysis indicates that the proposed methods lead to exceptional improvement in the order of complexity when implemented on GPU.
+This signifies a significant algorithmic advancement in solving matrix equations, with reduced computational cost and improved scalability.
+The presented methodology not only connects traditional numerical methods with neural architectures, but also contributes to interpretable [55, 56] and training-free network designs grounded in physical and mathematical principles.
+Future directions include the extension of this framework to nonlinear systems and time-dependent differential equations, and eigenvalue problems, enabling broader applications in scientific computing and model-driven learning.
+Code Availability
+A simple implementation of the proposed method and the code used to generate the data and the plots are available at the following link: https://github.com/nithingoona/HOJM